# Strong Haken via Thin Position

### v1

@@ 5. Amalgamation
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented generalized Heegaard surface $\mathcal{H}$. By Corollary 4.7, there exists a set of spheres $P\subset\mathcal{H}^{-}$ such that $M|_{P}$ is irreducible and so that $P$ contains a non-separating sphere $P_{0}$. Let $\mathcal{J}=\mathcal{H}\setminus P_{0}$ and observe it is an oriented acyclic generalized Heegaard surface for $M|_{P_{0}}$ with $\operatorname{net}\chi(\mathcal{J})=\operatorname{net}\chi(\mathcal{H})-2=2g(H)-4$. Consistently with a height function, amalgamate $\mathcal{J}$ to a Heegaard surface $J$ for $M|_{P_{0}}$ with $\operatorname{net}\chi(J)=\operatorname{net}\chi(\mathcal{J})$. Consequently, $g(J)=g(H)-1$. Thus, $g(M|_{P_{0}})\leq g(M)-1$. Since, $M|_{P_{0}}$ is homeomorphic to $M|_{P}$, we have $g(M|_{P})=g(M)-1$.
We now prove the additivity result. Assume that $M=M_{1}\#M_{2}$. By the previous result and uniqueness of prime decompositions of 3-manifolds, we may assume that none of $M$, $M_{1}$, $M_{2}$ contain a non-separating sphere.
@@
$22g(H)-2=\operatorname{net}\chi(H)=\operatorname{net}\chi(\mathcal{H})=2g(H_{1})+2g(H_{2})-$ 2 .
@@
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented $\mathcal{H}$. There exists a set of spheres $P\subset\mathcal{H}^{-}$ such that $M|_{P}$ is irreducible. By uniqueness of prime decompositions of 3-manifolds, we may group the components of $M|_{P}$ into those that are prime factors of $M_{1}$ and those that are prime factors of $M_{2}$. As above, by reconsituting $M_{1}$ and $M_{2}$ from their factors, we may assemble the components of $\mathcal{H}\setminus P$ into oriented, acyclic generalized Heegaard surfaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ for $M_{1}$ and $M_{2}$. (This may involve replacing components of $P$ with other thin spheres.) Since every sphere is separating, by the calculation at the start, we see that $g(M)\geq g(M_{1})+g(M_{2})$. Thus, equality holds. ∎
@@
If $H$ is a Heegaard surface for a reducible 3-manifold $M$, then $H$ is reducible.
@@
If $\partial M$ contains more than one sphere, it is easy to find an essential sphere intersecting $H$ in a single curve. If $\partial M$ consists of a single sphere, Haken’s Lemma follows from Haken’s Lemma for the manifold obtained by capping off the sphere with a 3-ball. Thus, for simplicity, we may assume that $\partial M$ does not contain a sphere. Let $\mathcal{H}$ be a thin generalized Heegaard surface obtained by thinning $H$ (Theorem 4.4). If $M$ is reducible, then by Corollary 4.7, $\mathcal{H}^{-}$ contains at least one essential sphere $P$. By Lemma 5.1 and Corollary 5.5, we may amalgamate $\mathcal{H}$ to a Heegaard surface $J$ such that $|J\cap P|=1$. By Theorem 5.10, $J$ is isotopic to $H$. That isotopy takes $P$ to a reducing sphere for $H$. ∎
@@
Figure 6. An example of spherical self-amalgamation of a Heegaard surface $H$ across a nonseparating sphere $P$. The compressionbody above $H$ in $M\setminus P$ is $A$ and the one below $H$ is $B$. The figure shows the two scars of $P$ in $M|_{P}$, one on either side of $H$. The Heegaard surface $H^{\prime}$ is formed by moving $H$ across one of the scars (by compressing along a semi-compressing disc) and then attaching a tube passing through $P$.
@@
Suppose that $M$ is a 3-manifold containing a nonseparating collection of spheres $P$ and that $H$ is an oriented Heegaard surface for $M\setminus P$ such that for each component $P_{0}\subset P$ the surface $H$ separates the two scars of $P_{0}$. Let $H^{\prime}\subset M$ be the surface constructed as follows. Let $A$ and $B$ be the compressionbodies of $M\setminus P$ above and below $H$ respectively. Let $\Delta_{A}\subset A$ be a collection of semi-compressing discs for $H$ such that $A|_{\Delta_{A}}$ is the disjoint union of $P\times I$ and a compressionbody equivalent to the result of capping off the components of $\partial_{-}A\cap P$ with 3-balls. Assume that $\Delta_{A}$ was chosen so that there is exactly one scar in each component of $P\times I$. Let $\alpha$ be a collection of arcs with endpoints in $H|_{\Delta_{A}}$ and interiors disjoint from $H|_{\Delta_{A}}$ chosen so that:
@@
    Each scar in $\Delta_{A}$ contains exactly one endpoint of $\alpha$
@@
As above, let $A$ and $B$ be the compressionbodies above and below $H$ in $M\setminus P$, respectively. Let $P\times[-1,1]$ be a regular neighborhood of $P$ in $M$ so that $P\times[-1,0]\subset B$ and $P\times[0,1]\subset A$. The $\partial$-reduction of $A$ along $\Delta_{A}$ results in compressionbodies embedded in $M$. The compressionbody equivalent to the result of capping of the components of $\partial_{-}A\cap P$ does not contain any sphere of $P$. Tubing $H$ along the arcs of $\alpha$ adds 1-handles to the positive boundary of that compressionbody and so it is a compressionbody $A^{\prime}$ embedded in $M$. Hence, $H^{\prime}$ is separating. The other side $B^{\prime}$ of $H^{\prime}$ can be constructed as follows. Begin by letting $B^{\prime}=B$. Let $\Delta_{B}$ be a complete set of compressing and semi-compressing discs for $B$, each disjoint from $\alpha$. The result of $\partial$-reducing $B$ along $\Delta_{B}$ is the union of 3-balls and product compressionbodies whose union contains $P\times[-1,0]$. When we we $\partial$-compress $A$ along $\Delta_{A}$, include $P\times[0,1]$ into $B^{\prime}$. Thus, $P\subset B^{\prime}$ and $B^{\prime}\setminus P$ is equivalent to $B$. The result of $\partial$-reducing $B^{\prime}$ along $\Delta_{B}$ is now the union of 3-balls, product compressionbodies, and components homeomorphic to $(P_{0}\times[-1,1])$, one for each component $P_{0}$ of $P$. Removing a regular neighborhood of $\alpha$ from $B^{\prime}$ to make it part of $A^{\prime}$, converts each of those into a $D^{2}\times[-1,1]$, which is a 3-ball. Undoing the $\partial$-reductions along $\Delta_{B}$, we see that $B^{\prime}$ is a compressionbody.
@@
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface containing a minimal complete collection of non-separating spheres. Then there is a Heegaard surface $H$ for $M$ amalgamation-obtained from $\mathcal{H}$. Furthermore, $H$ is unique up to isotopy.
@@
The existence of $H$ follows immediately from Lemmas 5.7 and 5.14 and Corollary 5.10. By Corollary 5.10, we need only show that the resulting Heegaard surface does not depend (up to isotopy) on the choice of $P$. Suppose that $P^{\prime}\subset\mathcal{H}^{-}$ is another minimal complete nonseparating collection of spheres. It is enough to consider the case when $P$ and $P^{\prime}$ differ by a single sphere. Let $P_{0}\subset P$ and $P^{\prime}_{0}\subset P^{\prime}$ be the sphere such that $P\setminus P_{0}=P^{\prime}\setminus P^{\prime}_{0}$. Let $M^{\prime}=M\setminus(P\cup P^{\prime}_{0})=M\setminus(P^{\prime}\cup P_{0})$. It has two components, and $\mathcal{H}$ restricts to an acyclic oriented generalized bridge surface in each. Choose a height function for each and amalgamate consistent with the height function to create Heegaard surfaces $H_{1}$ and $H_{2}$. Since $\mathcal{H}$ was oriented, both $H_{1}\cup P_{0}\cup H_{2}$ and $H_{1}\cup P^{\prime}_{0}\cup H_{2}$ are acyclic oriented generalized Heegaard surfaces for $M\setminus P^{\prime}$ and $M\setminus P$ respectively. Amalgamate these to Heegaard surfaces $J^{\prime}$ and $J$ respectively. Observe that each of $J^{\prime}$ and $J$ is isotopic to the result of connect-summing the Heegaard surfaces $H_{1}$ and $H_{2}$. As noted above, performing spherical self-amalmation of $J^{\prime}$ across $P^{\prime}$ is isotopic to the result of performing the connected sum of $J^{\prime}$ with $|P^{\prime}|$ copies of the genus 1 Heegard surface of $S^{2}\times S^{1}$. Similarly, the self-amalgamation of $J$ across $P$ is isotopic to the result of performing the connected sum of $J$ with $|P|$ copies of the genus 1 Heegaard surface of $S^{2}\times S^{1}$. Since $|P|=|P^{\prime}|$, the resulting Heegaard surfaces are isotopic. ∎

### v2

@@ 5. Amalgamation
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented generalized Heegaard surface $\mathcal{H}$. By Corollary 4.7, there exists a set of spheres $\mathcal{S}\subset\mathcal{H}^{-}$ such that $M|_{\mathcal{S}}$ is irreducible and so that $\mathcal{S}$ contains a nonseparating sphere $P_{0}$. Let $\mathcal{J}=\mathcal{H}\setminus P_{0}$ and observe it is an oriented acyclic generalized Heegaard surface for $M|_{P_{0}}$ with $\operatorname{net}\chi(\mathcal{J})=\operatorname{net}\chi(\mathcal{H})-2=2g(H)-4$. Consistently with a height function, amalgamate $\mathcal{J}$ to a Heegaard surface $J$ for $M|_{P_{0}}$ with $\operatorname{net}\chi(J)=\operatorname{net}\chi(\mathcal{J})$. Consequently, g(J)=g(H)1g(J)=g(H)-1. Thus, $g(M|_{P_{0}})\leq g(M)-1$. Since, $M|_{P_{0}}$ is homeomorphic to $M|_{P}$, we have $g(M|_{P})=g(M)-1$.
We now prove the additivity result. Assume that $M=M_{1}\#M_{2}$. By the previous result and uniqueness of prime decompositions of 3-manifolds, we may assume that none of $M$, $M_{1}$, $M_{2}$ contain a nonseparating sphere.
@@
$2g(H)-2=\operatorname{net}\chi(H)=\operatorname{net}\chi(\mathcal{H})=2g(H_{1})+2g(H_{2})-2.$
@@
Now suppose that $H$ is a Heegaard surface for $M$ of minimal genus. Thin $H$ to a locally thin acyclic oriented $\mathcal{H}$. There exists a set of spheres $\mathcal{S}\subset\mathcal{H}^{-}$ such that $M|_{\mathcal{S}}$ is irreducible. By uniqueness of prime decompositions of 3-manifolds, we may group the components of $M|_{\mathcal{S}}$ into those that are prime factors of $M_{1}$ and those that are prime factors of $M_{2}$. As above, by reconstituting $M_{1}$ and $M_{2}$ from their factors, we may assemble the components of $\mathcal{H}\setminus\mathcal{S}$ into oriented, acyclic generalized Heegaard surfaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ for $M_{1}$ and $M_{2}$. (This may involve replacing components of $\mathcal{S}$ with other thin spheres.) Since every sphere is separating, by the calculation at the start, we see that $g(M)\geq g(M_{1})+g(M_{2})$. Thus, equality holds. ∎
@@
Every Heegaard surface $H$ for a compact, connected, orientable reducible 3-manifold $M$ is reducible.
@@
As we mentioned in the introduction, if $\partial M$ contains a sphere, then every $\partial$-parallel sphere is a reducing sphere. We can easily isotope $H$ to intersect such a sphere in a single closed curve. (Likely the curve is inessential on $H$.) Hence, if $\partial M$ contains a sphere, $H$ is reducible. On the other hand, we can also show that if $M$ contains an essential sphere (whether or not $\partial M$ contains spheres), then there is an essential sphere in $M$ intersecting $H$ in a single simple closed curve. When $\partial M$ contains spheres, this is a marginally stronger version of Haken’s lemma than what we have stated.
Assume, therefore, that $M$ contains an essential sphere. In particular, it is not the connected sum of two or three 3-balls. Let $\widehat{M}$ be the result of capping off $\partial M$ with 3-balls. Any essential sphere in $\widehat{M}$ that is disjoint from the 3-balls we used to cap off $\partial M$ is also essential in $M$. If $\widehat{M}$ has no essential spheres, then $M$ is the connected sum of at least four 3-balls; in particular $\partial M$ contains at least four spheres. The Heegaard surface $H$ is also a Heegaard surface for $\widehat{M}$.
Suppose, first, that $M$ is the connected sum of at least four 3-balls. Let $P_{1},P_{2}$ be two spheres of $\partial M$ lying on the same side of $H$. Let $P\subset M\setminus H$ be a sphere separating $P_{1}$ and $P_{2}$ from both $H$ and all the other spheres in $\partial M$. Then $P$ is an essential sphere in $M$. We may isotope it so that it intersects $H$ in a single simple closed curve; likely the curve is inessential in $H$. Hence, there is an essential sphere in $M$ intersecting $H$ in a single closed curve, as claimed.
@@
It remains to consider the case when $\widehat{M}$ contains an essential sphere. Let $\mathcal{H}$ be a locally thin generalized Heegaard surface for $\mathcal{H}$ obtained by thinning $H$ (Theorem 4.4). By Corollary 4.7, $\mathcal{H}^{-}$ contains at least one essential (in $\widehat{M}$) sphere $P$. By Lemma 5.1 and Corollary 5.5, we may amalgamate $\mathcal{H}$ to a Heegaard surface $J$ such that $|J\cap P|=1$. By Theorem 5.10, $J$ is isotopic to $H$. That isotopy takes $P$ to a reducing sphere $P^{\prime}$ for $H$ in $\widehat{M}$. As $H$ is disjoint from the 3-balls $\widehat{M}\setminus M$, we may further isotope $P^{\prime}$, relative to $H$, to be disjoint from those 3-balls. The sphere $P^{\prime}$ is then an essential sphere in $M$ intersecting $H$ in a single simple closed curve, as desired. ∎
We can also prove the Structure Theorem for Weakly Reducible Heegaard surfaces.
Theorem 5.13 (Structure Theorem for Weakly Reducible Heegaard Surfaces).
Suppose that $H$ is an oriented weakly reducible Heegaard surface for a compact, orientable 3-manifold (possibly with boundary) $M$. Then there exists an oriented incompressible surface $F\subset M$ such that:
  1. (1)
    For each component $F^{\prime}$ of $F$, $F^{\prime}\cap H$ consists of a single simple closed curve, which is inessential in $F^{\prime}$. The curves of $F\cap H$ bound discs in $F$ all on the same side of $H$; we may perform the construction so that the discs lie on which ever side we wish.
  2. (2)
    $F\cap H$ separates $H$.
  3. (3)
    For each component $F^{\prime}$ of $F$, $\chi(F^{\prime})<-\chi(H)$.
  4. (4)
    There is a weak reducing pair of discs separated by $F$.
  5. (5)
    If some component of $F$ is a 2-sphere bounding a 3-ball, then there is a stabilizing pair of discs for $H$ disjoint from $F$
  6. (6)
    If some component of $F$ is $\partial$-parallel, then there is a $\partial$-stabilizing disc for $H$ disjoint from $F$.
  7. (7)
    There is a generalized Heegaard surface $\mathcal{H}$ for $M$ such that $\mathcal{H}^{+}$ is isotopic to $H+F$, the Haken sum of $H$ and $F$, and $\mathcal{H}^{-}=F$; furthermore, the thick surfaces $\mathcal{H}^{+}$ are strongly irreducible.
Proof.
The proof is much the same as our proof of Haken’s Lemma. Let $\mathcal{H}$ be a locally thin generalized Heegaard surface obtained by thinning a weakly reducible Heegaard surface $H$ (Theorem 4.4). Let $F=\mathcal{H}^{-}$ and note that $\mathcal{H}^{+}$ is strongly irreducible. As before, we may amalgamate $\mathcal{H}$ to a Heegaard surface $J$ such that for each component $F^{\prime}\subset\mathcal{H}^{-}$, we have $|F^{\prime}\cap J|=1$. By the construction, the curve $F^{\prime}\cap J$ is essential in $J$ and inessential in $F^{\prime}$. We note that in our process of amalgamation, we always extend the feet of 1-handles lying on the thick surface above a thin surface through the thin surface to lie on the thick surface below the thin surface. Consequently, each curve in $F\cap H$ bounds a disc in $F$ below $H$. We could instead have extended in the other direction, ensuring that each curve of $F\cap H$ bounds a disc in $F$ lying above $H$. This would be equivalent to reversing the orientation on $\mathcal{H}$, performing the amalgamation as we have described it, and then reverting the orientation back to the original. In any case, it is straightforward to confirm that the Haken sum $H+F$ of $H$ and $F$ is, after a small isotopy, disjoint from $F$ and is, in the complement of $F$, isotopic to $\mathcal{H}^{+}$ as claimed. This shows Conclusions (1) and (7).
In $M$, the dual digraph to $F$ can be constructed from the dual digraph to $\mathcal{H}$ by collapsing each edge corresponding to a thick surface to a single vertex. Thus, the dual digraph to $F$ is acyclic. Consequently, $F$ separates $M$ and, therefore, separates $J$. This is Conclusion (2).
If $F^{\prime}\subset F$ is a component, observe that it can be obtained from $H$ by a sequence of compressions, at least one of which is a compression along a compressing (rather than semi-compressing) disc. Thus, $\chi(F^{\prime})<-\chi(H)$. This is Conclusion (3).
Consider the construction of $J$. When we amalgamate, according to our construction, we extend a foot of a single 1-handle through each component of $F$. Let $D_{-}$ be the core disc for one of those 1-handles. In our construction, $D_{-}$ is an s-disc for a thick surface $H_{+}$ such that $D_{-}$ is in the compressionbody below $H_{+}$. The foot is extended so as to lie on a Heegaard surface $H_{-}$ in such a way that it misses the scars from a complete set of s-discs for the compressionbody above $H_{+}$. Let $D_{+}$ be one of those s-discs. Then $F$ separates $D_{-}$ from $D_{+}$. In subsequent amalgamations, although we do need to slide the feet of 1-handles onto other one handles to ensure that when we amalgamate each component of $F$ intersects $J$ in a single loop, we can do so without sliding the foot of a 1-handle across $F$; thus, discs $D_{-}$ and $D_{+}$ remain disjoint and separated by $F$ throughout the amalgamation process. This is Conclusion (4).
Suppose that some component $F^{\prime}$ of $F$ is a 2-sphere bounding a 3-ball in $M$. Since all closed surfaces in 3-balls are either spheres or are compressible, we may pass to an innermost such sphere. It bounds a 3-ball containing a unique component $H^{\prime}$ of $\mathcal{H}^{+}$. That component is not a 2-sphere, for then $H^{\prime}$ and $F$ could be amalgamated, contradicting the thinness of $\mathcal{H}$. Since $H^{\prime}$ is strongly irreducible, by Waldhausen’s classification of Heegaard surfaces of $S^{3}$, it is a torus and admits a stabilizing pair. When performing the amalgamations, we can ensure that the stabilizing pair remains disjoint from $F$. This is Conclusion (5).
If some component $F^{\prime}$ of $F$ is $\partial$-parallel, a similar argument shows the existence of a destabilizing disc disjoint from $F$. We use Scharlemann-Thompson’s classification of Heegaard splittings of $\text{(surface)}\times I$ in place of the Waldhausen classification. This is Conlcusion (6). ∎
@@
[…]
Figure 7. An example of spherical self-amalgamation of a Heegaard surface $H$ across a nonseparating sphere $P$. The compressionbody above $H$ in $M\setminus P$ is AA and the one below $H$ is $B$. The figure shows the two scars of $P$ in $M|_{P}$, one on either side of $H$. The Heegaard surface $H^{\prime}$ is formed by moving $H$ across one of the scars (by compressing along a semi-compressing disc) and then attaching a tube passing through $P$.
@@
Suppose that $M$ is a 3-manifold containing a nonseparating collection of mutually disjoint spheres $P$ and that $H$ is an oriented Heegaard surface for $M\setminus P$ such that for each component $P_{0}\subset P$ the surface $H$ separates the two scars of $P_{0}$. Let $H^{\prime}\subset M$ be the surface constructed as follows. Let $A$ and $B$ be the compressionbodies of $M\setminus P$ above and below $H$ respectively. Let $\Delta_{A}\subset A$ be a collection of semi-compressing discs for $H$ such that $A|_{\Delta_{A}}$ is the disjoint union of $P\times I$ and a compressionbody equivalent to the result of capping off the components of $\partial_{-}A\cap P$ with 3-balls. Assume that $\Delta_{A}$ was chosen so that there is exactly one scar in each component of $P\times I$. Let $\alpha$ be a collection of arcs with endpoints in $H|_{\Delta_{A}}$ and interiors disjoint from $H|_{\Delta_{A}}$ chosen so that:
@@
    Each scar from $\Delta_{A}$ contains exactly one endpoint of $\alpha$
@@
As above, let $A$ and $B$ be the compressionbodies above and below $H$ in $M\setminus P$, respectively. Let $P\times[-1,1]$ be a regular neighborhood of $P$ in $M$ so that $P\times[-1,0]\subset B$ and $P\times[0,1]\subset A$. The $\partial$-reduction of $A$ along $\Delta_{A}$ results in compressionbodies embedded in $M$. The compressionbody equivalent to the result of capping of the components of $\partial_{-}A\cap P$ does not contain any sphere of $P$. Tubing $H$ along the arcs of $\alpha$ adds 1-handles to the positive boundary of that compressionbody and so it is a compressionbody $A^{\prime}$ embedded in $M$. Hence, $H^{\prime}$ is separating. The other side $B^{\prime}$ of $H^{\prime}$ can be constructed as follows. Begin by letting $B^{\prime}=B$. Let $\Delta_{B}$ be a complete set of compressing and semi-compressing discs for $B$, each disjoint from $\alpha$. The result of $\partial$-reducing $B$ along $\Delta_{B}$ is the union of 3-balls and product compressionbodies whose union contains $P\times[-1,0]$. When we $\partial$-compress $A$ along $\Delta_{A}$, include $P\times[0,1]$ into $B^{\prime}$. Thus, $P\subset B^{\prime}$ and $B^{\prime}\setminus P$ is equivalent to $B$. The result of $\partial$-reducing $B^{\prime}$ along $\Delta_{B}$ is now the union of 3-balls, product compressionbodies, and components homeomorphic to $(P_{0}\times[-1,1])$, one for each component $P_{0}$ of $P$. Removing a regular neighborhood of $\alpha$ from $B^{\prime}$ to make it part of $A^{\prime}$, converts each of those into a $D^{2}\times[-1,1]$, which is a 3-ball. Undoing the $\partial$-reductions along $\Delta_{B}$, we see that $B^{\prime}$ is a compressionbody.
@@
Suppose that $\mathcal{H}$ is an oriented generalized Heegaard surface containing a minimal complete collection of nonseparating spheres. Then there is a Heegaard surface $H$ for $M$ amalgamation-obtained from $\mathcal{H}$. Furthermore, $H$ is unique up to isotopy.
@@
The surface $\mathcal{H}\setminus P$ is an oriented acyclic generalized Heegaard surface for $M\setminus P$. By Lemma 5.7, it admits a height function. By Corollary 5.10, $\mathcal{H}\setminus P$ may be amalgamated to an oriented Heegaard surface $J$ for $M\setminus P$ and $J$ is independent of the height function and the sequence of amalgamations. Note that $J\cup P$ is an oriented generalized Heegaard surface for $M$. Let $H$ be the result of spherical self-amalgamation of $J\cup P$ across $P$. By Lemma 5.15 it is a Heegaard surface for $M$ and it is amalgamation obtained from $\mathcal{H}$.
We need only show that $H$ does not depend (up to isotopy) on the choice of $P$. Suppose that $P^{\prime}\subset\mathcal{H}^{-}$ is another minimal complete nonseparating collection of spheres. It is enough to consider the case when $P$ and $P^{\prime}$ differ by a single sphere. Let $P_{0}\subset P$ and $P^{\prime}_{0}\subset P^{\prime}$ be the sphere such that $P\setminus P_{0}=P^{\prime}\setminus P^{\prime}_{0}$. Let $M^{\prime}=M\setminus(P\cup P^{\prime}_{0})=M\setminus(P^{\prime}\cup P_{0})$. It has two components, and $\mathcal{H}$ restricts to an acyclic oriented generalized Heegard surface in each. Choose a height function for each and amalgamate consistent with the height function to create Heegaard surfaces $H_{1}$ and $H_{2}$. Since $\mathcal{H}$ was oriented, both $H_{1}\cup P_{0}\cup H_{2}$ and $H_{1}\cup P^{\prime}_{0}\cup H_{2}$ are acyclic oriented generalized Heegaard surfaces for $M\setminus P^{\prime}$ and $M\setminus P$ respectively. Amalgamate these to Heegaard surfaces $J^{\prime}$ and $J$ respectively. Observe that each of $J^{\prime}$ and $J$ is isotopic to the result of connect-summing the Heegaard surfaces $H_{1}$ and $H_{2}$. As noted above, performing spherical self-amalmation of $J^{\prime}$ across $P^{\prime}$ is isotopic to the result of performing the connected sum of $J^{\prime}$ with $|P^{\prime}|$ copies of the genus 1 Heegard surface of $S^{2}\times S^{1}$. Similarly, the self-amalgamation of $J$ across $P$ is isotopic to the result of performing the connected sum of $J$ with $|P|$ copies of the genus 1 Heegaard surface of $S^{2}\times S^{1}$. Since $|P|=|P^{\prime}|$, the resulting Heegaard surfaces are isotopic. ∎